I am having trouble trying to gather how exactly to set up this problem. I tried mentally and drawing out what it could be asking if you could please help me.
George wants to use square pavers that are 6 inches on each side to completely surround his ower garden. The garden is 12 feet long and 3 feet wide. How many pavers should George buy?

Answer by Theo(13342)

George wants to use square pavers that are 6 inches on each side to completely surround his ower garden. The garden is 12 feet long and 3 feet wide. How many pavers should George buy?

convert the dimensions of the pavers to feet and you get each paver is 1/2 feet on each side.

this means that the length of each paver is 1/2 foot and the width of each paver is 1/2 foot.

to cover the length of the garden, you will need 12 / (1/2) = 24 pavers.

to cover the width of the garden, you will need 3 / (1/2) = 6 pavers.

you will need an additional 4 pavers to cover the corners.

the total number of pavers required is (24 * 2) + (6 * 2) + (1 * 4) = 64.

the garden will look like this.

24 tiles along the lenth on top and bottom.
6 tiles along the width on left and right side.
4 tiles on the corners.

$$$